[resolved] OIT OS and App deployment system (SCCM) is unavailable
- Update 12:00 – The issue with SCCM has been resolved and all SCCM functions should be operating normally
Issues with the SCCM system have appeared after updating the SCCM version this morning. OIT is working with Microsoft to resolve the issue. Currently OS deployment/App deployment/SCCM console may work intermittently. We will update this notification with any updates.
[RESOLVED] Helpdesk ticket creation and updates unavailable
** RESOLVED **
The issue with helpdesk ticket creation and updates have been resolved.
Creation of new helpdesk tickets or updates to existing tickets are currently unavailable. We anticipate this issue to be resolved in the morning. Email updates to tickets can still be sent and will be queued until the issue is resolved.
We apologize for the inconvenience
On 5/4 @ 2AM – several systems down for emergency maintenance
The following services may be unavailable for approximately 15 minutes starting at 2AM on 5/4 while we resolve an issue with the host supporting these systems:
University Police – AIM
Financial Aid site
OIT Labs Camera services
Help desk ticket updates via email
We apologize for the inconvenience,
[resolved] Automatic network drive mappings to shares on fauandi1 are intermittent
UPDATE: fauandi1 automatic drive mapping should now be working for most users following a logoff/logon. Some users may still be affected and we are working with the colleges to resolve.
The OIT-IPSEC-ENABLE GPO was unlinked from the top level OU last night as part of the andisec maintenance. This GPO specified the ‘Configure user Group Policy loopback processing mode’ as ‘Merge’ and it overrode any other GPO with that setting configured. We’re now seeing other colleges and departments with GPOs that specify that setting with ‘Replace’ which breaks the processing of the fauandi drive mapping GPO. We’re advising those colleges and departments to update their GPOs to ‘Merge’ to resolve the issue
We are receiving reports of some users not getting their fauandi1 shares mapped automatically. We are currently investigating the issue. As a workaround the shares can still be accessed by manually entering the share path in Windows Explorer.
[Resolved] Temporary disruption of DNS services
OIT will be restarting all DNS systems at 12:00PM. The system restarts will be staggered to minimize impact to computers on the FAU network. Please note you may experience slowed response times when contacting other systems over the network or internet as a result of DNS request time-outs. We expect all DNS systems to return fully to service by 12:30PM.
[resolved] DHCP is unavailable on Davie campus
-Update 12:56PM, DHCP service on Davie campus has returned to normal
We are experiencing an issue with the DHCP service on Davie campus. We are currently working to resolve the issue.
[resolved] OS deployment with SCCM is currently unavailable
After this morning’s SCCM update, OS deployment has stopped functioning. We are working to resolve the issue.
Thank you for your patience,
Intermittent Lync issues
[Update 1:08PM] – Microsoft has acknowledged the issue with their Lync service on 365. Microsoft engineers are currently working to resolve the issue.
Some users are reporting issues with Lync when signing-in. We are currently investigating this issue and its relation to the maintenance performed this morning.
[RESOLVED]Fort Lauderdale authentication and infrastructure support systems down
**Update 11:53AM – We removed a defective CPU from the system and all services are back and running on the remaining CPU. These services may run slightly slower on FTL campus until we can replace the defective CPU next week.
We are experiencing a hardware issue with the system that supports Windows authentication as well as DHCP and DNS at the Fort Lauderdale campus. These services at FTL are currently being supported by other campuses and we expect no end-user impact at this time. We are working to resolve this issue and will update when we have more information.
[resolved] Exchange mail unavailable for some users
Some users on Exchange were unable to access their email between Sunday morning and Monday morning. Network maintenance on Sunday morning prompted the migration of several mailboxes to the disaster recovery site. All mailboxes have been moved back to Boca and all email services are currently available. We are investigating why some networks are unable to contact the Exchange servers at the disaster recovery site.